The highest-energy electrons in Europium, Eu, are filling the __________.

A.

6f orbitals

B.

6d orbitals

C.

4f orbitals

D.

5f orbitals
Answer #1

Option C

Since electronic configuration of Eu= [Xe]4f76s2

Highest energy orbital=4f, (n+l value= 4+3=7, l=3 for f orbital)